### v3

# SaVe-Tag: Semantic-aware Vicinal Risk Minimization for Long-Tailed Text-Attributed Graphs

Abstract
Real-world graph data often follows long-tailed distributions, making it difficult for Graph Neural Networks (GNNs) to generalize well across both head and tail classes. Recent advances in Vicinal Risk Minimization (VRM) have shown promise in mitigating class imbalance with numeric interpolation; however, existing approaches largely rely on embedding-space arithmetic, which fails to capture the rich semantics inherent in text-attributed graphs. In this work, we propose our method SaVe-Tag (Semantic-aware Vicinal Risk Minimization for Long-Tailed Text-Attributed Graphs), a novel VRM framework that leverages Large Language Models (LLMs) to perform text-level interpolation, generating on-manifold, boundary-enriching synthetic samples for minority classes. To mitigate the risk of noisy generation, we introduce a confidence-based edge assignment mechanism that uses graph topology as a natural filter to ensure structural consistency. We provide theoretical justification for our method and conduct extensive experiments on benchmark datasets, showing that our approach consistently outperforms both numeric interpolation and prior long-tailed node classification baselines. Our results highlight the importance of integrating semantic and structural signals for balanced and effective learning on text-attributed graphs. The source code is available here: https://github.com/LWang-Laura/SaVe-TAG.
Figure 1: Comparison of numeric and LLM-based interpolation.(a) Illustration of the two interpolation strategies; (b) Advantages of LLM-based interpolation in producing boundary-enriching samples are highlighted (visualization uses synthetic data for illustration); (c) UMAP projection of the original and interpolated samples from the Cora dataset—the original data boundary is outlined, with numeric samples mostly staying within and LLM-generated samples extending beyond.
1 Introduction
Graph Neural Networks (GNNs) have proven effective for node classification by modeling structural dependencies in graph data. However, real-world graphs frequently exhibit long-tail distributions [1, 2, 3], making it difficult to learn from underrepresented classes.
Recent approaches to data imbalance often adopt the Vicinal Risk Minimization (VRM) principle [4], which minimizes loss over a synthetic neighborhood of the training data—extended from Empirical Risk Minimization (ERM) [5] that focuses solely on the existing data points. Typically, VRM is instantiated via interpolation techniques like Mixup [6] and SMOTE [7], and has been adapted for long-tailed graph learning in methods including GraphSMOTE [8] and MixupForGraph [9], etc. However, these methods often apply mathematic operations over input embeddings or hidden features, failing to preserve the rich manifold structure of text-attributed graphs (TAGs).
Figure 2: Effect of VRM.
Advances in large language models (LLMs) present new opportunities for data augmentation by offering external knowledge and creative generation. However, prior LLM-based augmentation approaches typically rely on zero-shot or few-shot prompting [10, 11], which usually fail to generate samples within the neighborhood of minority instances, as required by the VRM principle. Nevertheless, VRM has been demonstrated critical in improving long-tailed node classification. As shown in Figure 2a, VRM-guided numeric interpolation with nearest neighbors yields synthetic embeddings that outperform ERM-style embedding duplication.
Motivated by this insight, we hypothesize that prompting strategies explicitly designed to mimic interpolation will surpass standard zero/few-shot approaches, which is subsequently validated by the results in Figure 2b. Furthermore, integrating controlled prompting into LLM generation can alleviate common pitfalls of numeric interpolation, enabling manifold-aware, boundary-enriching augmentation that effectively exploits the knowledge encoded in LLMs (see Figure 1).
While LLM-based interpolation shows great potential, it also carries the risk of generating noisy samples. To mitigate this, we leverage the inherent structure of the graph as a natural filter. Using a confidence-based edge assignment strategy, noisy samples tend to remain isolated, limiting their influence on downstream tasks. In contrast, synthesized nodes aligned with the original distribution will be connected with the original graph, allowing GNN layers to reinforce their alignment through neighborhood aggregation.
In summary, we introduce a novel interpolation framework, SaVe-Tag, which unifies textual semantics and graph topology. Our method employs LLM-based interpolation to expand the vicinal boundary and reduce vicinal risk, followed by a topology-aware filtering mechanism that preserves alignment within the underlying data distribution. Our key contributions are summarized as follows:
  • To the best of our knowledge, we present the first semantic-aware VRM framework that employs LLM-based interpolation for node classification on long-tailed text-attributed graphs (TAGs).
  • We provide theoretical insights showing that LLM interpolation can generate manifold-preserving, boundary-enriching samples that effectively minimize vicinal risk, while a topology-aware filter helps mitigate noisy generation in the synthesized data.
  • We conduct extensive experiments demonstrating that our method outperforms embedding-based and long-tail graph baselines, highlighting the value of topology-aware filters and semantic benefits from LLMs.
2 Problem Formulation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$ be an attributed graph with $|\mathcal{V}|=n$ nodes, feature matrix [^1] For text-attributed graphs, each row of $\mathcal{X}$ stores a tokenized document. $\mathcal{X}\in\mathbb{R}^{n\times d}$, and labels $\mathcal{Y}=\{y_{v}|v\in\mathcal{V}\}$ with $y_{v}\in\!\{1,\dots,C\}$. Class imbalance induces a skewed empirical distribution $\mathcal{P}$ on $(\mathcal{X},\mathcal{Y})$. Our goal is to learn a node-level classifier $f_{\theta}:\mathbb{R}^{d}\!\!\times\!\mathcal{V}\!\to\!\Delta^{C-1}$, e.g., a GNN, that maximizes classification performance.
Figure 3: An overview of SaVe-Tag. Given an input graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$, we perform LLM-based interpolation to synthesize boundary-enriching samples $\hat{t}$ that minimize vicinal risk. A confidence function is then pre-trained on the original graph and later used to assign edges to the synthetic nodes via a top-$k$ selection strategy. When the resulting augmented graph $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$ is processed by a GNN, such edge assignment incorporates well-aligned nodes into their vicinity while isolating noisy samples.
3 SaVe-Tag Framework
An overview of our framework is shown in Figure 3, with details in Algorithm 1 from Appendix B.
Vicinal twins for tail classes.
Let $\mathcal{C}_{t}\subseteq\{1,\dots,C\}$ denote the set of tail (minority) classes. For each node $v_{1}\in\mathcal{V}$ with label $y_{v_{1}}\in\mathcal{C}_{t}$, we define its set of vicinal twins as
$\mathcal{N}_{k}^{\text{ (tail)}}(v_{1}):=\bigl{\{}v_{2}|\,v_{2}\in\mathcal{V}\;,\;y_{v_{2}}=y_{v_{1}},\;v_{2}\in\mathcal{N}_{k}(v_{1})\bigr{\}},$
i.e., the nodes from the same tail class that lie within the $k$-hop (or $k$-nearest) neighborhood $\mathcal{N}_{k}(v_{1})$ of $v_{1}$. As $(v_{1},v_{2})$ are in the vicinity, synthesizing new nodes from them follows the VRM principle.
LLM-based interpolation.
For any arbitrary pair of vicinal twins $(v_{i},v_{j})\in\mathcal{P}$, we query a temperature-controlled, class-conditioned LLM $\mathcal{L}_{\tau}$ to create $\hat{t}_{ij}\;=\;\mathcal{L}_{\tau}\!\bigl{(}t_{v_{i}},\,t_{v_{j}}\,\bigm{|}\,y_{v_{i}}\bigr{)}$ for $\hat{v}_{ij}$ such that $y_{\hat{v}_{ij}}=y_{v_{i}}$ and collect the new nodes in $\bar{\mathcal{V}}:=\{\hat{v}_{ij}\}$ with texts $\bar{\mathcal{T}}:=\{\hat{t}_{ij}\}$.
Edge Assignment with Confidence Function.
A frozen encoder $\phi:\mathcal{T}\!\to\!\mathbb{R}^{d^{\prime}}$ produces $z_{v}=\phi(t_{v}),\;z_{\hat{v}}=\phi(\hat{t})$, while a confidence function $\kappa:[0,1]$ is trained on the original graph. Each $\hat{v}$ connects to the $k$ highest-scoring neighbors
$\mathcal{N}_{k}(\hat{v})=\operatorname*{arg\,top}_{k,\;u\in\mathcal{V}}\bigl{[}\kappa_{\omega}(z_{u})\,\langle z_{\hat{v}},z_{u}\rangle\bigr{]},\qquad\bar{\mathcal{E}}_{\hat{v}}=\{(\hat{v},u)\mid u\in\mathcal{N}_{k}(\hat{v})\},$
and is identified as isolated if $\max_{u}\kappa_{\omega}(z_{u})<\tau$. Setting $\mathcal{V}^{\prime}=\mathcal{V}\cup\bar{\mathcal{V}},\;\mathcal{E}^{\prime}=\mathcal{E}\cup\!\!~{}\bigcup_{\hat{v}}\bar{\mathcal{E}}_{\hat{v}},\;\mathcal{T}^{\prime}=\mathcal{T}\cup\bar{\mathcal{T}}$ yields $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$, on which the GNN $f_{\theta}$ is further trained.
4 SaVe-Tag Theory
4.1 LLM-based Interpolation
In this subsection, we formalize how class-conditioned generation preserves semantic consistency, enriches boundary regions, and improves generalization through Vicinal Risk Minimization.
4.1.1 Setup and Notation
Embeddings and class manifolds.
Let $\phi:\mathcal{X}\to\mathbb{R}^{d^{\prime}}\!,~{}z_{v}\;:=\;\phi(x_{v})$ denote the penultimate embedding produced by $f_{\theta}$. For each class $c$ we write $\mathcal{M}_{c}=\bigl{\{}z_{v}:y_{v}=c\bigr{\}}$ for the (empirical) class manifold in feature space and set $\mathcal{M}:=\bigcup_{c=1}^{C}\mathcal{M}_{c}$.
Text generator.
We use $\mathcal{L}_{\tau}$ to denote the LLM text generator with temperature $\tau$. Given seeds $t_{1},t_{2}\!\in\!\mathcal{T}$ (raw texts) and an explicit class prompt $c$, the generator returns $\hat{t}\;=\;\mathcal{L}_{\tau}(t_{1},t_{2}\mid c).\;$
4.1.2 Manifold-preserving Interpolation
Theorem 4.1 (Off-Manifold Numeric Interpolation).
Let $\phi(t_{1}),\phi(t_{2})\in\mathcal{M}_{c}$. Under the linearity constraint, numeric interpolation yields the combination based on convex assumption [12]:
$z_{\lambda}\;=\;\lambda\,\phi(t_{1})\;+\;(1-\lambda)\,\phi(t_{2})~{}\text{ for}~{}\lambda\in(0,1)$
If the class manifold $\mathcal{M}_{c}$ is non-convex, then ${}~{}z_{\lambda}\notin\mathcal{M}_{c}~{}~{}\text{ for all }\lambda\in(0,1)~{}$ [13, 14].
Theorem 4.2 (Manifold-Preserving Class-Consistent Generation).
Let $t_{1},t_{2}\in\mathcal{T}$ satisfy ${}\phi(t_{1}),\;\phi(t_{2})\;\in\;\mathcal{M}_{c}~{}$, and $\hat{t}=\mathcal{L}_{\tau}\bigl{(}t_{1},\,t_{2}\mid c\bigr{)}$ condition on class prompt $c$
Then there exists a small $\delta>0$ such that, with probability at least  $1-\delta$ , $\phi(\hat{t})\;\in\;\mathcal{M}_{c}~{},$
i.e. the LLM preserves class consistency and manifold under label-guided generation:
[…]
Proof.
By conditioning the model on the target class $c$, prompting transforms $\mathcal{L}_{\tau}$ into a class-conditional generator: $\hat{t}\sim p\bigl{(}t\mid t_{1},t_{2},c\bigr{)}.$ If the model has learned the conditional support of $\mathcal{M}_{c}$ accurately—e.g., through ample in-class examples—then with probability at least $1-\delta$, the sampled $\hat{t}$ will lie within the semantic manifold of class $c$, i.e. $\phi(\hat{t})\in\mathcal{M}_{c}$. ∎
4.1.3 Boundary-enriching Effect of LLM Samples
Intuitively, the greatest gain from augmentation arises when synthetic points fill sparsely-populated decision-boundary zones [15]. We therefore quantify how many generated texts land in these border regions and later show (Theorem 4.6) that a higher fraction yields a lower vicinal risk.
Let $\gamma(x)=f_{\theta}^{(y)}(x)-\max_{j\neq y}f_{\theta}^{(j)}(x)$ be the logit margin; write $\gamma_{\min}(S)=\min_{x\in S}\gamma(x)$ for a set $S$ [16]. Denote the augmented training set by $\tilde{\mathcal{D}}=\{(x_{i},y_{i})\}_{i=1}^{n}\cup\{(\hat{x}_{j},\hat{y}_{j})\}_{j=1}^{m}.$
Definition 4.3 (Boundary-Coverage Rate (BCR)).
Following [17], a synthetic example $\hat{x}$ is a boundary sample (a borderline) iff the $k$-NN majority label on $\mathcal{M}$ differs from $\hat{y}$. BCR is the fraction of such samples in $\tilde{\mathcal{D}}$.
Theorem 4.4 (Margin lower bound).
Fix $\delta>0$. Let $f_{\theta}^{\mathrm{orig}}$ be trained on $\mathcal{D}$ with margin $\gamma_{\min}(\mathcal{D})=\gamma_{0}$. Augment with $m$ LLM-generated points and retrain to obtain $f_{\theta}^{\mathrm{aug}}$. Assume each boundary sample satisfies $\lvert\gamma(\hat{x}_{j})\rvert\leq\delta$. Then, adapting the proof of [18],
$\gamma_{\min}\!\bigl{(}\tilde{\mathcal{D}}\bigr{)}\;\;\geq\;\;\gamma_{0}-\delta\bigl{(}1-\mathrm{BCR}\bigr{)}.$
Hence more boundary samples $\Rightarrow$ a tighter realized margin.
Next, we show that such boundary-rich distribution can also minimize vicinal risk.
4.1.4 Vicinal Risk Minimization with an LLM Neighborhood
Vicinal Risk Minimization (VRM) defines a local surrogate loss around each training point. When the neighborhood is populated with boundary samples, this surrogate becomes a sharper proxy for the true generalization error. We formalize the link below.
Theorem 4.5 (On-manifold Vicinal Risk Theorem).
Let $\{(t_{i},c_{i})\}_{i=1}^{n}$ be the training set and define the LLM-based vicinal distribution
$\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})=\sum_{j=1}^{n}\alpha_{ij}\,\delta_{\,\hat{t}=\mathcal{L}_{\tau}(t_{i},t_{j}\mid c_{i})},\qquad\sum_{j}\alpha_{ij}=1.$
The corresponding vicinal risk is $R_{vrm}(f_{\theta})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\hat{t}\sim\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})}\!\bigl{[}\mathbf{L}\bigl{(}f_{\theta}(\phi(\hat{t})),c_{i}\bigr{)}\bigr{]}.$
If  Theorem 4.2 holds, $R_{vrm}(f_{\theta})$ is evaluated on-manifold with probability $\geq 1-\delta$.
Theorem 4.6 (Boundary-coverage $\Rightarrow$ vicinal-risk reduction).
Let $f_{\theta}^{\mathrm{aug}}$ be obtained from VRM training on $\tilde{\mathcal{D}}$. Denote by $L$ the Lipschitz constant of loss $\mathbf{L}$ in its first argument. Then, for any $\eta>0$,
$R_{vrm}(f_{\theta}^{\mathrm{aug}})\;\;\leq\;\;R_{vrm}(f_{\theta}^{\mathrm{orig}})\;-\;L\,\gamma_{0}\,\bigl{(}\mathrm{BCR}-\eta\bigr{)}\;+\;\mathcal{O}(\delta).$
In words: increasing BCR lowers the vicinal risk up to constants dictated by the original margin and the loss smoothness. Hence the boundary-enriching effect of LLM samples (Sec. 4.1.3) directly translates to a tighter VRM bound. See Appendix F for detailed proofs.
4.2 Graph as Natural Filters
We now show that our edge assignment mechanism with confidence function naturally acts as a filter: well-aligned synthetic nodes are pulled toward the correct class manifold, whereas noisy nodes are automatically isolated. Throughout this section we retain the notation introduced in Section 4.1.1.
4.2.1 Setup and Notation
Let $\bar{\mathcal{V}}$ be the set of synthetic nodes with textual attributes $\bar{\mathcal{T}}$ after interpolation. Define $\mathcal{V}^{\prime}=\mathcal{V}\cup\bar{\mathcal{V}},~{}\mathcal{T}^{\prime}=\mathcal{T}\cup\bar{\mathcal{T}}.$ We then form the augmented graph $\mathcal{G}^{\prime}=\bigl{(}\mathcal{V}^{\prime},\,\mathcal{E}^{\prime},\,\mathcal{T}^{\prime}\bigr{)},$ where $\mathcal{E}^{\prime}\setminus\mathcal{E}$ is obtained by evaluating, for each synthetic node $\hat{v}\in\bar{\mathcal{V}}$ and each node $u\in\mathcal{V}^{\prime}$, whether to include the edge $(\hat{v},u)$.
Aggregator.
Following [19, 20], any message-passing GNN layer is abstracted as
$h^{(\ell+1)}_{v}\;=\;\alpha\,\mathbf{W}\,h^{(\ell)}_{v}\;+\;(1-\alpha)\!\!\sum_{u\in\mathcal{N}(v)}\!\beta_{vu}\,\mathbf{W}\,h^{(\ell)}_{u},$ (1)
where $h^{(\ell)}_{v}\!\in\!\mathbb{R}^{p}$, $\alpha\!\in\![0,1]$, $\beta_{vu}\!\geq\!0$ and $\sum_{u\in\mathcal{N}(v)}\beta_{vu}=1$.
Representation manifolds.
After training, each class $c$ occupies a roughly low-dimensional manifold $\mathcal{M}_{c}\!\subset\!\mathbb{R}^{p}$ in representation space [21]. Following [22], we write $\text{dist}(x,\mathcal{S}):=\inf_{y\in\mathcal{S}}\|x-y\|_{2}$.
4.2.2 Confidence–based Edge Function
Definition 4.7 (Confidence function).
A measurable map $\kappa:\mathbb{R}^{d^{\prime}}\!\times\!\mathbb{R}^{d^{\prime}}\!\to\![0,1]$ assigns a confidence score to any pair of text embeddings. Given a threshold $\tau\!\in\!(0,1)$ (or a global top–$k$ rule), we add $(u,v)$ to $\mathcal{E}^{\prime}$ iff $\kappa(z_{u},z_{v})\geq\tau$.
4.2.3 Confidence $\Rightarrow$ Pulling
Theorem 4.8 (Pulling well-aligned nodes).
Let $\hat{v}$ be a synthetic node labelled $c$ and it has a nonempty set of neighbors $\mathcal{N}^{\prime}(\hat{v})=\{u\in V:\kappa(z_{\hat{v}},z_{u})\geq\tau\}\neq\varnothing$. Assume every neighbor $u\in\mathcal{N}^{\prime}(\hat{v})$ satisfies $\text{\emph{dist}}\bigl{(}h^{(\ell)}_{u},\mathcal{M}_{c}\bigr{)}\leq\varepsilon$. Based on  [23, 24], there exists $\gamma\!\in\!(0,1)$, independent of $\kappa$, such that
$\text{\emph{dist}}\bigl{(}h^{(\ell+1)}_{\hat{v}},\mathcal{M}_{c}\bigr{)}\;\leq\;\gamma\;\text{\emph{dist}}\bigl{(}h^{(\ell)}_{\hat{v}},\mathcal{M}_{c}\bigr{)}.$
Iterating the Equation (1) yields exponential convergence of $h^{(\ell)}_{\hat{v}}$ toward $\mathcal{M}_{c}$  [25].
4.2.4 No confidence $\Rightarrow$ Isolation
Theorem 4.9 (Isolating noisy nodes).
If $\kappa(z_{\hat{v}},z_{u})<\tau$ for every $u\!\in\!V$ then $\mathcal{N}^{\prime}(\hat{v})=\varnothing$ and $h^{(\ell+1)}_{\hat{v}}=\alpha\,\mathbf{W}\,h^{(\ell)}_{\hat{v}}\,$ for all layers $\ell$. Consequently, $\hat{v}$ neither influences nor is influenced by $\mathcal{G}^{\prime}$.
Proof.
Immediately from Definition 4.7 and Equation 1. ∎
5 Empirical Results
In this section, we present extensive experiments to validate our proposed method, emphasizing its benefits from both Semantic and Structural perspectives. We further evaluate the integrated framework to demonstrate its effectiveness in addressing node classification on long-tailed graphs. Specifically, our experiments aim to answer the following three research questions (RQs):
  • RQ1: Can large language models (LLMs) effectively generate boundary-enriching samples with manifold interpolation, enabling VRM for textual data distributions?
  • RQ2: Can our confidence-based edge assignment effectively filter out noisy samples?
  • RQ3: Does the entire proposed framework outperform existing state-of-the-art methods in graph learning on long-tailed TAGs?
We first describe our experimental setup, followed by an in-depth analysis to address the stated research questions. Our results consistently show that the proposed method significantly outperforms existing baselines, signifying its effectiveness for node classification on long-tailed TAGs.
5.1 Experimental Setup
5.1.1 Datasets
Following prior studies on long-tailed graph learning [26, 27, 9], we use Cora [28], PubMed [29], and Citeseer [29], along with three larger Amazon datasets, namely Photo, Computer, and Children [30], for scalability analysis.
[…]
5.1.2 Metrics
Following prior work on long-tailed graph classification [3, 2, 26, 27, 9, 8], we evaluate model performance using four standard metrics: balanced accuracy (bAcc), macro-F1 score (F1), geometric mean (GMean), and traditional accuracy (Acc). All experiments are repeated five times with different fixed seeds, and we report the mean performance across these runs.
To further assess boundary-enriching effects, we report the Boundary Coverage Ratio (BCR), as defined in Definition 4.3, and the Boundary Proximity Score (BPS). BPS is computed as the reciprocal of the Trust Score [31], measuring the ratio between the distance to the nearest out-of-class centroid and the distance to the in-class centroid. Additionally, we report the In-Class Rate (ICR), defined as the percentage of samples predicted with the same label as their ground truth. Implementation details for these metrics are provided in Section 5.2.
5.1.3 Baselines
We compare our method against two categories of baselines: (1) LLM-based data augmentation, including zero shot and few shot approaches [10, 11]; and (2) long-tailed graph learning, including Oversampling [32], SMOTE [7], Embed-SMOTE[33], MixupForGraph [9], GraphSMOTET[8], GraphSMOTEO[8], LTE4G[3], and HierTail [2]. See Appendix B for details.
To comprehensively evaluate our method, we design and implement three variants of SaVe-Tag, each corresponding to a specific VRM-based strategy. Detailed prompting strategies for each variant are provided in the Appendix B.
  • SaVe-TagO: Mirrors the oversampling method but generate a texts similar to the existing one rather than simple duplication from the conventional method.
  • SaVe-TagS: Inspired by SMOTE [7], interpolates with nearest neighbors within the same class.
  • SaVe-TagM: Extends SMOTE [7] by interpolating embeddings with nearest neighbors from potentially different classes, inspired by Mixup techniques.
5.1.4 Implementation Details
By default, we generate texts with Llama3-8B-Instruct and use Sentence-BERT (all-MiniLM-L6-v2) as the text encoder $\phi(\cdot)$. Node classification is performed using a Graph Convolutional Network (GCN). See Appendix B for details about hyperparameters.
5.2 Semantic-aware Benefits (RQ1)
On-Manifold. To compare on-manifold LLM-based and off-manifold numeric interpolation, we evaluate node classification on multiple benchmarks. Figure 4(a) shows results on Citeseer. Embed leverages LLaMA 3.2-1B embeddings for interpolation, while SimCSE and SBERT serve as alternative embedding-based baselines. Our LLM-based method using LLaMA 3.2-1B consistently outperforms all numeric interpolation baselines. Similar trends on other datasets can be found in Appendix A.
Figure 4: LLM interpolation preserves manifolds and enriches boundaries on the Citeseer dataset.
Method Cora Pubmed Citeseer Acc F1 GMean bAcc Acc F1 GMean bAcc Acc F1 GMean bAcc Original 72.5$\pm$1.4 70.4$\pm$1.1 84.5$\pm$0.6 74.7$\pm$0.9 55.1$\pm$8.2 47.8$\pm$15.0 62.2$\pm$7.9 51.5$\pm$9.7 53.7$\pm$0.1 49.6$\pm$2.2 71.4$\pm$0.9 55.9$\pm$1.3 Zero-shot 68.3$\pm$0.5 63.3$\pm$0.7 81.1$\pm$0.4 69.4$\pm$0.7 71.4$\pm$0.4 71.4$\pm$0.5 78.3$\pm$0.4 72.3$\pm$0.5 64.8$\pm$0.5 62.1$\pm$0.5 78.0$\pm$0.3 65.2$\pm$0.5 Few-shots 73.9$\pm$0.6 72.7$\pm$0.6 85.4$\pm$0.2 76.1$\pm$0.3 72.0$\pm$0.3 72.0$\pm$0.3 79.2$\pm$0.2 73.6$\pm$0.3 68.2$\pm$0.9 65.6$\pm$0.7 79.8$\pm$0.4 67.8$\pm$0.5 LLMO 73.7$\pm$0.7 72.1$\pm$0.7 85.3$\pm$0.3 76.1$\pm$0.4 71.5$\pm$0.3 71.4$\pm$0.3 78.8$\pm$0.3 73.1$\pm$0.4 69.6$\pm$0.8 67.0$\pm$0.6 80.5$\pm$0.4 68.8$\pm$0.5 LLMM 74.2$\pm$0.4 72.6$\pm$0.3 85.6$\pm$0.1 76.5$\pm$0.2 71.9$\pm$0.4 72.0$\pm$0.4 79.0$\pm$0.2 73.4$\pm$0.3 69.5$\pm$0.6 66.8$\pm$0.4 80.3$\pm$0.2 68.6$\pm$0.3 LLMS 74.5$\pm$0.4 72.5$\pm$0.5 85.5$\pm$0.2 76.3$\pm$0.3 72.5$\pm$0.4 72.5$\pm$0.4 79.4$\pm$0.3 73.8$\pm$0.3 69.2$\pm$0.6 66.5$\pm$0.5 80.1$\pm$0.3 68.3$\pm$0.5
Table 1: Comparison of VRM-based LLM interpolation with standard LLM augmentation. Subscripts O, M, and S denote the prompting strategies of their corresponding SaVe-Tag variants, but without edge assignment. Best and second-best results are emboldened and underlined, respectively.
Boundary-enriching Effects. We first visualize the embeddings from LLM-based and numeric interpolation. Figure 4(b) shows the UMAP projection of the Citeseer dataset, with the boundary of the original data circled. While numeric interpolated samples tend to stay within the original boundary, LLM-generated samples often extend beyond it, enlarging the VRM decision region.
We continue the evaluation with BCR and BPS metrics, as shown in the Figure 4c and the Appendix A. These metrics are computed for three settings: the original data, the combined set of original and interpolated data (Interp+Orig), and the interpolated data alone (Interp-Only). The results indicate that the original data typically achieves the highest BCR and BPS values, suggesting that real-world data tends to lie near decision boundaries. Notably, our LLM-based interpolation consistently outperforms its numerically interpolated counterpart in both metrics, better aligned with the original distribution with enriched boundary regions. .
In addition, we train a separate MLP classifier for each dataset to evaluate the class consistency of the augmented data. Each classifier is trained on a balanced dataset, where all classes have an equal number of training samples. Figure 4(d) shows the predictions on augmented samples from Citeseer, where the classifier achieves an accuracy of $71.3\pm 0.7$. As shown, numerically interpolated samples achieve 100% In-Class Rate (ICR) for all augmented classes, indicating complete class consistency. In contrast, LLM-interpolated samples yield ICRs of 81%, 69%, 94%, and 100%, suggesting that some synthesized samples lie near class boundaries and may be predicted as neighboring classes. Additional results on other datasets are provided in the Appendix A.
LLM as Vicinal Risk Minimization. We evaluate the effectiveness of LLM-based interpolation within the Vicinal Risk Minimization (VRM) framework. As shown in Table 1, interpolation-based prompting significantly enhances the quality of synthesized textual data, resulting in improved performance on downstream node classification tasks.
5.3 Necessity of Confidence-based Edge Assignment (RQ2)
Isolating Noisy Samples.
Figure 5: Top-k edge selection.
We instantiate Definition 4.7 using a top–$k$ rule, where each synthetic node retains its $k$ highest-confidence edges. The value of $k$ is set to #$\#augmented\_nodes\times n$, and we vary $n$ to study its effect on node isolation and classification performance. Figure 5a shows that the number of isolated nodes ($\log$-scaled) drop as $\log(n)$ increases. Figure 5b reveals a bell-shaped accuracy curve with respect to $\log(n)$, measured relative to $n{=}1$. These trends support our hypothesis:
  • Too small $k$rightarrow isolation. Many synthetic nodes become singletons (Theorem 4.9), so the minority manifold is under-populated and performance suffers.
  • Too large $k$rightarrow over-smoothing. Excessive connectivity leads to repeated averaging in Equation (1) and pushing node representations towards a global mean, resulting in over-smoothing [34].
Effectiveness of Confidence Function.
We compare two edge assignment strategies for augmented nodes: edge duplication from the target node (without the ‘${C}$’ subscript) and edge assignment with our confidence-based function (denoted by ‘${C}$’). Table 2 shows that our confidence-based edge assignment consistently outperforms naive edge duplication under LLM-based interpolation. A similar trend holds for numeric interpolation, with the exception of Cora and Photo, likely due to the generation of off-manifold samples that our confidence mechanism successfully isolates.
Dataset Method SaVe-TagS SaVe-TagM SaVe-TagO F1 $\uparrow$ (%) ACC $\uparrow$ (%) F1 $\uparrow$ (%) ACC $\uparrow$ (%) F1 $\uparrow$ (%) ACC $\uparrow$ (%) Cora Origin 70.40 + 0.0 72.46 + 0.0 70.40 + 0.0 72.46 + 0.0 70.40 + 0.0 72.46 + 0.0 Num 70.59 + 0.1 73.11 + 0.9 73.21 + 3.9 75.32 + 3.8 71.04 + 0.9 73.80 + 1.9 NumC 70.09 - 0.4 72.44 - 0.9 71.86 + 2.0 71.18 - 1.8 70.89 + 0.7 72.76 + 0.4 LLM 72.54 + 3.0 74.51 + 2.8 72.30 + 2.7 74.21 + 2.4 70.09 - 0.4 72.86 + 0.6 LLMC 74.05 + 5.2 76.36 + 5.4 70.54 + 0.2 72.51 + 0.1 73.21 + 4.0 75.32 + 4.0 PubMed Origin 47.79 + 0.0 55.09 + 0.0 47.79 + 0.0 55.09 + 0.0 47.79 + 0.0 55.09 + 0.0 Num 69.06 + 44.5 69.43 + 26.0 70.48 + 47.5 70.73 + 28.4 69.44 + 45.3 69.77 + 26.6 NumC 69.19 + 44.8 69.63 + 26.4 72.66 + 52.0 72.46 + 31.5 69.01 + 44.4 69.43 + 26.0 LLM 72.54 + 51.8 72.50 + 31.6 71.34 + 49.3 71.38 + 29.6 71.44 + 49.5 71.51 + 29.8 LLMC 76.43 + 59.6 76.35 + 38.6 76.18 + 59.4 76.26 + 38.4 72.96 + 52.7 72.80 + 32.1 Citeseer Origin 49.63 + 0.0 53.71 + 0.0 49.63 + 0.0 53.71 + 0.0 49.63 + 0.0 53.71 + 0.0 Num 62.69 + 26.3 65.89 + 22.7 63.52 + 28.0 67.24 + 25.2 63.12 + 27.2 66.28 + 23.4 NumC 65.19 + 31.4 68.79 + 28.1 64.10 + 29.2 68.89 + 28.3 64.75 + 30.5 68.24 + 27.1 LLM 66.51 + 34.0 69.19 + 28.8 66.93 + 34.9 69.69 + 29.8 66.99 + 35.0 69.58 + 29.5 LLMC 66.85 + 34.7 69.89 + 30.1 67.96 + 36.9 71.60 + 33.3 66.82 + 34.6 69.90 + 30.1 Photo Origin 54.99 + 0.0 51.69 + 0.0 54.99 + 0.0 51.69 + 0.0 54.99 + 0.0 51.69 + 0.0 Num 57.29 + 4.2 55.74 + 7.8 57.92 + 5.3 56.04 + 8.4 57.08 + 3.8 55.08 + 6.6 NumC 56.05 + 1.9 54.97 + 6.3 56.74 + 3.2 54.57 + 5.6 54.44 - 1.0 51.31 - 0.7 LLM 58.40 + 6.2 56.75 + 9.8 59.53 + 8.3 58.06 + 12.3 59.32 + 7.9 58.19 + 12.6 LLMC 59.20 + 7.7 58.65 + 13.5 59.45 + 8.1 58.60 + 13.4 63.62 + 15.7 66.17 + 28.0 Computer Origin 44.93 + 0.0 49.34 + 0.0 44.93 + 0.0 49.34 + 0.0 44.93 + 0.0 49.34 + 0.0 Num 48.90 + 8.8 54.47 + 10.4 49.47 + 10.1 53.85 + 9.1 49.47 + 10.1 55.78 + 13.0 NumC 49.31 + 9.7 54.69 + 10.8 53.78 + 19.7 60.03 + 21.7 48.85 + 8.7 54.57 + 10.6 LLM 52.81 + 17.5 60.17 + 21.9 52.45 + 16.7 55.77 + 13.0 52.45 + 16.7 59.45 + 20.5 LLMC 56.89 + 26.6 66.26 + 34.3 57.83 + 28.7 66.98 + 35.8 57.03 + 26.9 65.12 + 32.1 Children Origin 2.25 + 0.0 3.00 + 0.0 2.25 + 0.0 3.00 + 0.0 2.25 + 0.0 3.00 + 0.0 Num 15.87 + 605.3 16.01 + 433.7 15.93 + 608.0 16.68 + 456.0 15.75 + 600.0 15.80 + 426.7 NumC 19.32 + 758.7 18.87 + 529.0 19.07 + 747.6 18.10 + 503.3 18.13 + 705.8 17.50 + 483.3 LLM 17.94 + 697.3 17.82 + 494.0 18.82 + 736.4 19.98 + 566.0 18.06 + 702.7 18.31 + 510.3 LLMC 19.69 + 775.1 18.64 + 521.3 22.16 + 884.9 24.50 + 716.7 21.81 + 869.3 22.99 + 666.3
Table 2: Ablation Studies. Node classification under three variants of SaVe-TAG. Origin uses no augmentation; Num and LLM apply numeric and LLM-based interpolation, respectively. Subscript C indicates the use of confidence-based edge generation versus naive edge duplication otherwise.
5.4 Enhancement in Long-tailed Graph Learning (RQ3)
Figure 6: Node Classification Improvement.
Performance.
Table 3 shows that our method, particularly SaVe-TagS, outperforms prior baselines in node classification on long-tailed text-attributed graphs.
Moreover, our method effectively narrows the disparity between the head (majority) and tail (minority) classes. As shown in Figure 6a, across all benchmarks, SaVe-TagS achieves a smaller accuracy gap than numeric interpolation, underscoring its effectiveness in mitigating class imbalance in long-tailed learning.
Stability.
We further perform a sensitivity analysis to evaluate the stability of SaVe-Tag  under varying imbalance ratios. As shown in Figure 6b, SaVe-TagS exhibits smaller fluctuations compared to other baselines as imbalance ratios vary. Additional results are shown in the Appendix  A.
Method Cora Pubmed Citeseer Acc F1 GMean bAcc Acc F1 GMean bAcc Acc F1 GMean bAcc SMOTE 71.8$\pm$2.7 70.8$\pm$2.1 82.7$\pm$1.7 71.8$\pm$2.7 68.5$\pm$6.2 65.5$\pm$10.1 75.9$\pm$4.8 68.5$\pm$6.2 50.7$\pm$3.0 47.9$\pm$3.1 67.6$\pm$2.2 50.7$\pm$3.0 Oversampling 72.9$\pm$1.3 72.0$\pm$0.9 83.4$\pm$0.9 72.9$\pm$1.3 58.5$\pm$6.3 50.4$\pm$6.5 68.1$\pm$5.1 58.5$\pm$6.3 52.0$\pm$4.9 49.6$\pm$4.7 68.5$\pm$3.6 52.0$\pm$4.9 EmbedSMOTE 70.0$\pm$1.6 69.8$\pm$1.4 81.5$\pm$1.0 70.0$\pm$1.6 63.4$\pm$7.6 60.3$\pm$9.7 72.0$\pm$5.9 63.4$\pm$7.6 47.5$\pm$2.0 45.4$\pm$2.5 65.2$\pm$1.5 47.5$\pm$2.0 GraphSMOTET 74.4$\pm$4.0 74.2$\pm$4.3 84.4$\pm$2.6 74.4$\pm$4.0 67.9$\pm$5.0 67.8$\pm$5.0 75.5$\pm$3.9 67.9$\pm$5.0 50.5$\pm$2.8 48.1$\pm$3.2 67.4$\pm$2.1 50.5$\pm$2.8 GraphSMOTEO 74.3$\pm$3.7 74.1$\pm$4.9 84.3$\pm$3.0 74.3$\pm$4.7 69.7$\pm$5.0 69.9$\pm$4.8 76.9$\pm$3.9 69.7$\pm$5.0 52.0$\pm$1.6 49.8$\pm$1.8 68.6$\pm$1.2 52.0$\pm$1.6 MixupForGraph 66.4$\pm$2.7 63.32$\pm$2.7 80.1$\pm$1.6 68.0$\pm$2.5 69.6$\pm$0.8 68.7$\pm$1.1 75.9$\pm$1.1 68.7$\pm$1.6 63.2$\pm$1.9 59.4$\pm$1.9 75.1$\pm$1.8 60.9$\pm$2.5 HierTail  [^1] Modified code while running on PubMed dataset to prevent out-of-memory issues due to scalability. 72.8$\pm$5.3 73.3$\pm$4.7 83.3$\pm$3.4 72.8$\pm$5.3 72.4$\pm$8.2 72.4$\pm$7.7 79.0$\pm$6.3 72.5$\pm$8.0 49.6$\pm$9.6 44.5$\pm$11.3 66.5$\pm$7.4 49.6$\pm$9.6 LTE4G  [^1] Modified code while running on PubMed dataset to prevent out-of-memory issues due to scalability. 74.2$\pm$3.8 74.1$\pm$3.8 84.3$\pm$2.4 74.2$\pm$3.8 72.9$\pm$1.5 72.6$\pm$1.9 79.4$\pm$1.2 72.9$\pm$1.5 51.4$\pm$3.8 49.2$\pm$3.7 68.1$\pm$2.8 51.4$\pm$3.8 SaVe-TagO 72.1$\pm$0.1 70.9$\pm$0.1 84.5$\pm$0.3 74.8$\pm$0.4 72.8$\pm$0.3 73.0$\pm$0.4 80.2$\pm$0.2 75.0$\pm$0.2 69.9$\pm$0.2 66.8$\pm$0.2 79.9$\pm$0.1 67.8$\pm$0.2 SaVe-TagM 73.0$\pm$0.4 71.8$\pm$0.3 85.1$\pm$0.2 75.8$\pm$0.4 71.3$\pm$0.2 71.2$\pm$0.3 79.1$\pm$0.1 73.6$\pm$0.3 69.5$\pm$0.3 65.9$\pm$0.4 79.3$\pm$0.3 66.9$\pm$0.5 SaVe-TagS 76.4$\pm$0.3 74.1$\pm$0.3 85.8$\pm$0.2 76.6$\pm$0.3 76.3$\pm$0.7 76.4$\pm$0.8 82.5$\pm$0.5 77.6$\pm$0.7 69.9$\pm$0.3 66.9$\pm$0.3 80.1$\pm$0.3 68.1$\pm$0.4
Table 3: Comparison of SaVe-Tag with previous long-tail graph learning baselines.
5.5 Ablation Studies
We conduct the ablation studies to show the significance of both LLM-based interpolation and confidence-based edge assignment, shown in Table 2.
LLM v.s. Num. Leveraging the generative capabilities of LLMs for VRM proves to be more effective than embedding interpolation. For instance, SaVe-TagS achieves a 10.3% increase in average accuracy on the PubMed dataset compared to its numeric counterpart.
w/ C, v.s. w/o C. With confidence function (denoted as C) enhances the node classification performance. Same analysis has been discussed in Section 5.3
6 Discussion
Related Work.
Here we outline relevant concepts and advances; see Appendix D for details.
  • Vicinal Risk Minimization (VRM): VRM [35] extends ERM [5] by estimating risk from local vicinity distributions, enhancing model generalization. It underlies methods like Mixup [12] and SMOTE [7], which create synthetic samples via interpolation.
  • Long-tailed Graph Learning: In graph learning, long-tailed node classification is typically addressed via hierarchical task grouping like HierTail [2] and expert models like LTE4G [3].
  • Large Language Model (LLM) Data Augmentation: Generally, LLMs augment data by generating contextually relevant text via zero-shot or few-shot instruction-based prompts [36, 10].
Limitations.
While LLM-based interpolation offers clear benefits, it is more time- and resource-intensive than numeric embedding methods. Accordingly, we provide a cost analysis in Appendix C. Despite the overhead, we find the trade-off worthwhile: our framework supports small, open-source language models, requires no extensive training or tuning, and applies LLMs only once during augmentation—allowing reuse across tasks. Most importantly, our method consistently outperforms baselines, validating the effectiveness of LLM interpolation in practice.
Another limitation is our exclusive focus on GNN backbones, overlooking non-graph alternatives such as MLPs and transformer-based encoders. This focus enabled a deeper exploration of topology-aware filtering, but limits the generalizability of our findings to non-graph models. To partially address this, Appendix A reports results with an MLP, which show similar trends and suggest that our method may generalize beyond GNNs or graph-structured data.
7 Conclusion
This paper presents a novel interpolation framework for long-tailed node classification that unifies semantic and structural signals. With LLMs, we generate manifold-consistent, boundary-enriching samples to extend VRM to the textual domain. To mitigate out-of-distribution generation, we introduce a confidence-based edge assignment that filters synthetic nodes via graph connectivity. Theoretically and empirically, our method outperforms numeric interpolation, standard LLM augmentation, and prior long-tailed graph learning across benchmarks.
Appendix A Additional experimental results
A.1 Semantic-aware Benefits.
UMAP Visualization.
Figure 7 shows the UMAP visualization across all six datasets, with the boundary of the original data circled. As we can see, for all the datasets, numeric interpolated samples tend to stay within the original boundary, while LLM-generated samples often extend beyond it, enlarging the VRM decision region.
Figure 7: UMAP projection of the original and interpolated samples across various benchmarks—the original data boundary is outlined, with numeric samples mostly staying within and LLM-generated samples extending beyond.
Boundary Coverage Rate (BCR) and Boundary Proximity Score (BPS).
Figure 8: Boundary Coverage Rate (BCR) and Boundary Proximity Score (BPS) for numeric (blue) vs. LLM (yellow) interpolation in Interp+Orig and Interp-Only settings across six datasets; the red dashed line marks the original baseline.
Figure 8 presents the evaluation results using BCR and BPS metrics across all datasets. We report performance under three settings: original data, the combination of original and interpolated samples (Interp+Orig), and interpolated samples alone (Interp-Only).
Across all benchmarks, our LLM-based interpolation consistently outperforms numeric interpolation on both metrics, suggesting that LLM-generated samples align more effectively with decision boundaries. Particularly in some cases, LLM-based interpolation even exceeds the original data in the BCR, highlighting its potential to extend and enrich the original decision boundary.
Figure 9: Node Classification Improvement.
In-Class Rate (ICR).
We train a separate MLP classifier on each dataset using a balanced set of class samples to assess class consistency. The average In-Class Rate (ICR)—the proportion of augmented samples classified into their intended class—is reported in Figure 9.
Numerical interpolation achieves 100% ICR on Cora, PubMed, and Citeseer, and nearly 100% on others, reflecting high class consistency. In contrast, LLM-based samples show lower ICRs, indicating proximity to decision boundaries that causes occasional misclassification by MLP.
Node classification Performance
We evaluate the impact of LLM-based and numeric interpolation on node classification across multiple benchmarks. Results are presented in Figure 10. Embed utilizes LLaMA 3.2-1B embeddings for interpolation, while SimCSE and SBERT serve as embedding-based numeric baselines. Across all datasets, the our LLM-based approach with LLaMA 3.2-1B consistently outperforms all numeric interpolation baselines.
Figure 10: Node classification performance (F1) under various implementation of SaVe-TagS. Embed uses Llama3.2-1B embeddings for interpolation. SimCSE and SBERT are compared as alternative embedding baselines. Our LLM-based method with Llama3.2-1B consistently outperforms.
A.2 Necessity of Topology-aware Filters
Figure 11: Impact of average assigned edges on performance for Photo (a), Children (b), and Computer (c). We report changes from the baseline ($n{=}1$) across four metrics: $\Delta$Acc, $\Delta$F1, $\Delta$bAcc, and $\Delta$GMean.
We instantiate Definition 4.7 using a top–$k$ rule, where each synthetic node retains its $k$ highest-confidence edges. The value of $k$ is set to #$\#augmented\_nodes\times n$, and we vary $n$ to study its effect on node isolation and classification performance. Figure 11 reveals a bell-shaped performance curve with respect to $\log(n)$ on Photo, Computer, and Children, measured relative to $n{=}1$. More discussion can be found in Section 5.3.
A.3 Enhancement in Long-tailed Graph Learning.
Stability.
We also adjusted the imbalance ratio, selecting from $\{0.1,0.2,0.3,0.4,0.5\}$, to examine how our model performs under different levels of imbalance. As shown in Figure 12, SaVe-TagS (the bold line) exhibits less performance drop as the imbalance ratio decreases, demonstrating less fluctuation compared to other baselines.
Figure 12: F1 score (%) versus imbalance ratio on (a) Cora, (b) PubMed, and (c) Citeseer, comparing other baselines of long-tailed graph learning and with our SaVe-TagS (the bold line).
Classifier Variants.
To assess the generalization of our method across different classifier architectures, we conduct additional ablation studies using one GNN-based backbone (SAGE) and one non-graph-based model (MLP). Since MLP does not incorporate structural information (i.e., it lacks neighborhood aggregation), the confidence-based edge assignment has no effect. Therefore, we report only the results without the subscript C. As shown in Table 4, SaVe-Tag  effectively addresses long-tailed classification with both GNNs and non-graph-based models. Its strong performance with MLP highlights its potential for text classification and broader applications.
Model Method Cora PubMed Citeseer Photo Computer Children F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc SaVe-TagS MLP Origin 50.31 52.34 25.51 42.59 61.70 64.64 4.25 7.48 3.09 4.61 1.06 2.03 Num 58.28 61.35 64.63 68.62 61.70 64.64 34.96 34.74 28.19 31.77 15.31 14.85 LLM 64.13 66.80 72.77 72.07 64.04 66.25 43.26 40.79 37.65 40.26 16.64 16.83 SAGE Origin 67.5 70.2 60.7 62.4 52.9 55.5 47.4 44.2 36.5 38.0 2.5 4.0 Num 70.6 73.1 68.3 68.6 65.0 68.0 54.2 52.1 43.1 47.3 18.7 18.6 NumC 70.3 72.4 69.2 69.6 65.8 69.4 50.6 48.9 49.3 54.7 19.4 18.7 LLM 72.5 74.5 72.5 72.5 67.4 70.0 57.7 56.0 48.6 53.7 20.1 20.8 LLMC 74.0 76.4 76.4 76.3 66.6 69.3 57.6 56.1 52.0 59.1 20.3 20.3 SaVe-TagO MLP Origin 50.31 52.34 25.51 42.59 61.70 64.64 4.25 7.48 3.09 4.61 1.06 2.03 Num 57.85 60.97 66.06 66.46 60.23 63.38 35.47 35.01 28.65 32.32 15.89 15.32 LLM 63.02 64.12 70.25 69.88 64.68 67.07 43.34 42.15 38.18 42.74 18.65 20.38 SAGE Origin 67.5 70.2 60.7 62.4 52.9 55.5 47.4 44.2 36.5 38.0 2.5 4.0 Num 70.6 73.8 69.0 69.3 65.3 68.6 54.6 53.5 43.0 47.1 18.5 19.0 NumC 69.1 70.9 68.3 68.6 65.5 68.9 49.0 46.1 42.9 46.0 18.6 19.1 LLM 72.1 73.7 71.4 71.6 67.5 70.3 57.8 56.1 48.1 53.9 20.9 21.8 LLMC 69.2 70.1 72.5 72.3 66.2 69.0 58.0 57.3 52.5 59.9 22.5 23.8 SaVe-TagM MLP Origin 50.31 52.34 25.51 42.59 61.70 64.64 4.25 7.48 3.09 4.61 1.06 2.03 Num 61.66 62.89 67.88 67.77 61.06 63.57 36.98 35.39 29.25 33.24 15.95 15.95 LLM 62.94 64.47 71.12 70.56 64.52 66.64 44.84 42.28 38.47 41.47 18.80 20.09 SAGE Origin 67.5 70.2 60.7 62.4 52.9 55.5 47.4 44.2 36.5 38.0 2.5 4.0 Num 71.9 74.0 69.8 70.0 64.8 68.7 56.0 54.1 46.0 51.4 19.4 19.8 NumC 69.8 70.1 69.9 69.7 64.3 68.6 53.6 51.7 46.6 51.4 19.2 19.6 LLM 72.3 74.2 71.5 71.3 68.0 70.8 58.4 56.7 49.1 54.8 21.7 23.1 LLMC 70.5 72.5 75.7 75.2 67.8 70.9 58.6 57.8 50.6 57.4 23.9 26.5
Table 4: Ablation Studies. Node classification performance using SAGE and MLP under three variants of SaVe-Tag. Origin denotes no augmentation; Num and LLM apply numeric and LLM-based interpolation, respectively. The subscript C indicates the use of confidence-based edge generation, while its absence denotes naive edge duplication. The best results for both MLP and SAGE are emboldened, and the second-best results for SAGE are underlined.
Appendix B Implementation Details
B.1 Code Implementation
Algorithm
Our codes are provided in the following link:
https://github.com/LWang-Laura/SaVe-TAG
The pseudocode of SaVe-Tag  is provided in Algorithm 1.
1
2
Input: Attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{T},\mathcal{E})$ with texts $\mathcal{X}$ and labels $\mathcal{Y}$;
vicinal order $k$; temperature–controlled LLM $\mathcal{L}_{\tau}$;
frozen encoder $\phi$; confidence network $\kappa_{\omega}$;
edge budget $k_{\text{edge}}$; confidence threshold $\tau$
Output: Augmented graph $\mathcal{G}^{\prime}=(\mathcal{V}^{\prime},\mathcal{T}^{\prime},\mathcal{E}^{\prime})$; trained classifier $f_{\theta}$
3 Initialize $\bar{\mathcal{T}}$, $\bar{\mathcal{V}}$, $\bar{\mathcal{E}}$ as empty sets
4 Identify tail classes: $\mathcal{C}_{t}\leftarrow\{c\mid\text{freq}(c)<\text{median freq}\}$
5 $\mathcal{V}_{t}\leftarrow\{v\in\mathcal{V}\mid y_{v}\in\mathcal{C}_{t}\}$
6
7
8 Function DataAugment($v,u$):
9       $\hat{t}\leftarrow\mathcal{L}_{\tau}\bigl{(}t_{v},t_{u}\,\big{|}\,y_{v}\bigr{)}$
10       Initialize new node $\hat{v}$ with text $\hat{t}$ and label $y_{\hat{v}}=y_{v}$
11       return $\hat{v}$
12
13
// Generate vicinal twins and interpolate with the LLM
14 for $v\in\mathcal{V}_{t}$ do
15       $\mathcal{N}^{(\mathrm{tail})}_{k}(v)\leftarrow\{u\in\mathcal{V}_{t}\mid u\in\mathcal{N}_{k}(v)\}$
16       for $u\in\mathcal{N}^{(\mathrm{tail})}_{k}(v)$ do
17             $\hat{v}\leftarrow\texttt{DataAugment}(v,u)$
18             Append $\hat{v}$ to $\bar{\mathcal{V}}$, its text $\hat{t}$ to $\bar{\mathcal{T}}$
19
20
21
22
// Pre-train confidence function on original graph
23 $\kappa_{\omega}\leftarrow\text{train}(\{z_{v}=\phi(t_{v})\}_{v\in\mathcal{V}})$
24
25
// Edge assignment guided by confidence
26 for $\hat{v}\in\bar{\mathcal{V}}$ do
27       $z_{\hat{v}}\leftarrow\phi(\hat{t}_{\hat{v}})$
28       $s_{u}\leftarrow\kappa_{\omega}(z_{u})\langle z_{\hat{v}},z_{u}\rangle$ for $u\in\mathcal{V}$
29       $\mathcal{N}_{k_{\text{edge}}}(\hat{v})\leftarrow\arg\mathrm{top}_{k_{\text{edge}}}s_{u}$
30       if $\max_{u}s_{u}\geq\tau$ then
31             ¯¯$\bar{\mathcal{E}}\leftarrow\bar{\mathcal{E}}\cup\{(\hat{v},u)\mid u\in\mathcal{N}_{k_{\text{edge}}}(\hat{v})\}$
32
33      else
34             pass;
35
36
37
38Form augmented graph:
39 $\mathcal{V}^{\prime}\leftarrow\mathcal{V}\cup\bar{\mathcal{V}}$,
40 $\mathcal{T}^{\prime}\leftarrow\mathcal{T}\cup\bar{\mathcal{T}}$,
41 $\mathcal{E}^{\prime}\leftarrow\mathcal{E}\cup\bar{\mathcal{E}}$
42
43Train GNN $f_{\theta}$ on $\mathcal{G}^{\prime}$
44
Result: $\mathcal{G}^{\prime},\,f_{\theta}$
45
Algorithm 1 SaVe-Tag
Computing Environment and Resources
We implement our model under PyG [37] and Sentence Transformer [38] modules. Experiments are conducted on a NVIDIA GeForce RTX 3090 and the OS was Ubuntu 22.04.4 LTS with 128GB RAM.
B.2 Hyperparameters
The detailed hyperparameters of our setups for reproducibility are listed below:
  • Text Encoding: We utilize the Sentence Transformer (all-MiniLM-L6-v2) [38], for its efficiency in semantic comprehension. This model strikes an excellent balance between being lightweight and delivering strong performance. A comparison with other embeddings is provided in Figure 10.
  • Node Classification: By default, we use GCN as our model, consisting of 2 hidden layers with 64 neurons each. The dropout rate is 0.5 and the model is trained for 1000 epochs with a learning rate of 0.01 (same configuration as we switch to SAGE when conducting experiments in Table 4).
  • Confidence Function: We employ MLP models composed of 1 hidden layer with 256 neurons and set with a dropout rate of 0, which are trained for 1000 epochs with a learning rate of 0.001.
  • Text Generation: By default, we leverage pre-trained Llama3-8B-Instruct from Meta for text generation, configured with bfloat16 and default parameters.
  • Data Augmentation: Due to the limited number of training nodes in our low-labeled, imbalanced settings, we selected $k=3$ for the k-nearest neighbors in our LLM-based methods.
  • Edge generation: For small datesets (Cora, PubMed, and Citeseer), we set $k=|\mathcal{V}|\times 20$ and add the top $k$ highest prediction score edges to the graph. For larger datasets (Photo, Computer, and Children), we increase $k$ to $|\mathcal{V}|\times 40$ correspondingly.
B.3 Datasets
In this part, we include brief descriptions of each dataset. For Cora and PubMed, we downloaded the preprocessed datasets from https://github.com/CurryTang/Graph-LLM [39]. As to Citeseer, Photo, Children, and Computer datasets, the preprocessed graph data are downloaded from https://github.com/CurryTang/TSGFM [40] and their detailed descriptions can be found in https://github.com/sktsherlock/TAG-Benchmark [30]. The category names for each dataset are listed below, with their indices corresponding to the numeric labels in the dataset.
  • Cora: [‘Rule Learning’, ‘Neural Networks’, ‘Case Based’, ‘Genetic Algorithms’, ‘Theory’, ‘Reinforcement Learning’, ‘Probabilistic Methods’]
  • PubMed: [‘Diabetes Mellitus, Experimental’, ‘Diabetes Mellitus Type 1’, ‘Diabetes Mellitus Type 2’]
  • Citeseer:[‘Agents’, ‘ML(Machine Learning)’, ‘IR(Information retrieval)’, ’DB(Database)’, ‘HCI (Human-computer Interaction)’, ‘AI(Artificial Intelligence)’]
  • Photo:[ ‘Video Surveillance’,"Accessories’,‘Binoculars & Scopes’, ‘Video’, ‘Lighting & Studio’, ‘Bags & Cases’,‘Tripods & Monopods’, ‘Flashes’, ‘Digital Cameras’, ‘Film Photography’, ‘Lenses’, ‘Underwater Photography’]
  • Children: [‘Literature & Fiction’, ‘Animals’, ‘Growing Up & Facts of Life’, ‘Humor’, ‘Cars Trains & Things That Go’, ‘Fairy Tales Folk Tales & Myths’, ‘Activities Crafts & Games’, ‘Science Fiction & Fantasy’, ‘Classics’, ‘Mysteries & Detectives’, ‘Action & Adventure’, ‘Geography & Cultures’, ‘Education & Reference’, ‘Arts Music & Photography’, ‘Holidays & Celebrations’, ‘Science Nature & How It Works’, ‘Early Learning’, ‘Biographies’, ‘History’, “Children’s Cookbooks", ‘Religions’, ‘Sports & Outdoors’, ’Comics & Graphic Novels’, ’Computers & Technology’]
  • Computer: [‘Computer Accessories & Peripherals’, ‘Tablet Accessories’, ‘Laptop Accessories’, ‘Computers & Tablets’, ‘Computer Components’, ‘Data Storage’, ‘Networking Products’, ‘Monitors’, ‘Servers’, ‘Tablet Replacement Part’]
B.4 Baselines Details
LLM-based Data Augmentation Baselines.
  • Zero-shot [10, 11]: Utilizing a pre-trained LLM to generate textual attributes for a given target label without providing specific examples.
  • Few-shot [10, 11]: Prompting an LLM to generate textual attributes for a given target label, providing examples for in-context learning.
Long-tailed Graph Learning Baselines.
  • Oversampling [32]: Repeatedly duplicates node embeddings that belong to the minority classes.
  • SMOTE [7]: Generating synthetic nodes by interpolating tail nodes with neighbors and assigning edges by copying their neighbors’ edges.
  • Embed-SMOTE [33]: Applying SMOTE on hidden layer embeddings rather than the input features.
  • MixupForGraph [9]: Synthesizing additional data for minority classes by interpolating randomly paired hidden representation as well as their labels
  • GraphSMOTE [8]: Interpolating minority nodes with their nearest neighbors and generating new edges using a co-trained link predictor, with two variants denoted by subscript T and O, depending on whether the predicted edges are discrete or continuous.
  • LTE4G [3]: Clustering nodes by class and degree, training subset-specific experts, and distilling them into head and tail students for classification.
  • HierTail [2]: Extracting shared class information via hierarchical task grouping and balancing head–tail gradient contributions.
SaVe-Tag Variants.
The description of the three variants (SaVe-TagS, SaVe-TagM, SaVe-TagO) are listed in Section 5.1.3. A case study illustration is provided in Figure 13.
Figure 13: A case study illustrating three variants our LLM-based interpolation: (a) SaVe-TagO, (b) SaVe-TagS, and (c) SaVe-TagM.
System: You are a helpful AI assistant for generating {Task} from {Dataset}, where each {Text} follows the format <START>{Format}<END>.
User: Give me the first {Text} from {Dataset} with topic [$\mathcal{C}_{1}$].
Assistant: <START>[$t_{1}$]<END>.
User: Give me the second {Text} from {Dataset} with topic [$\mathcal{C}_{2}$].
Assistant: <START>[$t_{2}$]<END>.
User: Give me the third {Text} from {Dataset} with topic [$\mathcal{C}_{2}$]. It should be more similar to the first {Text} and less similar to the second {Text}.
Assistant:
Table 5: Prompt template for SaVe-TagS and SaVe-TagM (${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{1}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{2}}$ for SaVe-TagO).
System: You are a helpful AI assistant for generating {Task} from {Dataset}, where each {Task} follows the format <START>{Format}<END>.
User: Give me the first {Task} from {Dataset} with topic [$\mathcal{C}_{1}$].
Assistant: <START>[$t_{1}$]<END>.
User: Give me the second {Task} from {Dataset} with topic [$\mathcal{C}_{1}$]. It should be more similar to the first {Task}.
Assistant:
Table 6: Prompt template for SaVe-TagO across all datasets.
[…]
Table 7: Input parameters for different datasets in the prompt templates.
B.5 Prompt Design
The prompt template for SaVe-TagO is shown in Table 6, while SaVe-TagS and SaVe-TagM share a common template in Table 5, with the constraint of shared class ( ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{1}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}_{2}}$ ) when using SaVe-TagS. Both templates are dataset-specific, with their parameters detailed in Table 7.
Appendix C Efficiency and Scalability Analysis
Name # Nodes # Edges # Class # Tail Ave Text Len # Batch Prompts/Batch Time/Prompt Cora 2,708 10,858 7 5 891 20 4 3.06s PubMed 19,717 88,670 3 2 1649 8 4 4.51s Citeseer 3186 8450 6 4 1022 16 4 3.50s Photo 48,362 500,939 12 8 804 64 2 4.60s Computer 87,229 721,081 10 6 499 24 2 4.09s Children 76,875 1,554,578 24 15 1255 240 1 7.21s
Table 8: Dataset Details. We recorded the following statistics on all datasets: the number of nodes, edges, and classes; the number of tail (minority) classes; the average text length of node attributes; the total number of batches; the number of prompts per batch (batch size); and the average time per prompt (generated entry) during generation.
C.1 Time Costs
For scalability analysis, we report the statistics regarding the size of the graph (number of edges, nodes, average length of text attributes, etc.) in the Table 8.
To facilitate text generation, we generate samples in batches. While larger batches reduce overall runtime, memory constraints limit us to small batch sizes. We then compute the average LLM inference time per generated sample. The data regarding batch size and inference time per prompt are also reported in Table 8.
C.2 Adaptability for Small Language Models
Figure 14: The F1 score and generation time per entry of LLM variants across different datasets implementing SaVe-TagS.
Figure 14 shows the performance and efficiency of SaVe-TagS using language models with different sizes, including Llama-3-8B-Instruct, Llama-3.2-1B-Instruct, and Llama-3.2-3B-Instruct. As expected, smaller models require less inference time for text generation. Notably, the F1 score remains relatively stable across model sizes, highlighting the adaptability of our method to lightweight, open-source LLMs—a promising direction for reducing generation costs.
Appendix D Related Works
Vicinal Risk Minimization (VRM)
VRM [35] is a principle in statistical learning theory that extends the traditional Empirical Risk Minimization (ERM) framework [5]. Introduced by Chapelle et al., VRM proposes estimating the expected risk by considering a vicinity distribution around each training sample, rather than relying solely on the empirical distribution. This approach aims to improve generalization by incorporating local neighborhoods of data points into the learning process. VRM has been foundational in developing techniques like Mixup [6] and SMOTE [7], which generate synthetic training examples through linear interpolations of existing data points, thereby enhancing model generalization capabilities.
  • Mixup [6]: Given $(x_{i},y_{i}),(x_{j},y_{j})$, draw $\lambda\sim\mathrm{Beta}(\alpha,\alpha)$ and set
    $\tilde{x}=\lambda x_{i}+(1-\lambda)x_{j},\quad\tilde{y}=\lambda y_{i}+(1-\lambda)y_{j}.$
  • SMOTE [7]: For minority sample $x_{i}$, choose a neighbor $x_{k}$, draw $\lambda\sim U(0,1)$ and define
    $\tilde{x}=x_{i}+\lambda\,(x_{k}-x_{i}),\quad\tilde{y}=y_{i}=y_{j}.$
Long-tail Graph Learning
In many real-world graph-structured datasets, the distribution of node classes follows a long-tailed pattern, where a few classes (head classes) have abundant samples, while many others (tail classes) have scarce representations. This imbalance poses significant challenges for graph neural networks (GNNs), which may become biased towards
  • Hierarchical Task Grouping: Formulating long-tailed classification as a multi-task learning problem, where each class prediction is treated as a separate task. This approach allows for shared representations among related tasks, improving performance on tail classes [2].
  • Expert Models: Training specialized models (experts) for different subsets of classes or nodes, such as the LTE4G [3], which considers both class and degree long-tailedness in graphs.
LLM Data Augmentation
Large Language Models (LLMs) have revolutionized data augmentation strategies, especially in natural language processing tasks. By generating synthetic data that mirrors the distribution and characteristics of real data, LLMs can enhance model training in low-resource settings. Key approaches include:
  • Instruction-based Augmentation: Empowering LLMs with task-specific instructions to generate diverse and contextually relevant data [36].
  • Zero-shot or Few-shot Augmentation: Enabling LLMs to generate synthetic examples with minimal or no task-specific labels via in-context learning[10].
Appendix E Ethics and Broader Impact
SaVe-Tag  LLMs to generate synthetic text for augmenting long-tailed node classification on text-attributed graphs. By mitigating class imbalance, it improves classification performance on widely used benchmark datasets such as Cora, Citeseer, and PubMed. The use of lightweight, open-source LLMs and one-time data augmentation helps reduce computational cost, potentially enabling broader access to graph learning techniques in low-resource settings.
Our study is conducted entirely on publicly available, anonymous datasets, and does not involve any personally identifiable or sensitive information. The generated synthetic data is stored in files and used exclusively for research purposes. While our method does not raise significant ethical concerns in its current scope, we encourage practitioners to clearly label synthetic data in downstream applications to avoid confusion between real and generated content.
Appendix F Detailed Proof
F.1 Proof of Theorem 4.4 (Margin lower bound)
Notation recap.
$\gamma(x)$ is the logit margin of a point $x$, $\gamma_{\min}(S)=\min_{x\in S}\gamma(x)$, and $\gamma_{0}:=\gamma_{\min}(\mathcal{D})$ is the smallest margin on the original training set $\mathcal{D}$. After augmentation, we have $\tilde{\mathcal{D}}\;=\;\mathcal{D}\cup\widehat{\mathcal{D}}_{\mathrm{bd}}\cup\widehat{\mathcal{D}}_{\mathrm{in}},$ where $\widehat{\mathcal{D}}_{\mathrm{bd}}$ (resp. $\widehat{\mathcal{D}}_{\mathrm{in}}$) holds all synthetic boundary (resp. interior) samples and $\mathrm{BCR}=\frac{|\widehat{\mathcal{D}}_{\mathrm{bd}}|}{|\widehat{\mathcal{D}}_{\mathrm{bd}}|+|\widehat{\mathcal{D}}_{\mathrm{in}}|}\in[0,1].$
Assumptions for this theorem.
  1. A1.
    Margin monotonicity. Retraining on $\tilde{\mathcal{D}}$ does not decrease any margin that was already $\geq\gamma_{0}$. Thus, every point in $\mathcal{D}\cup\widehat{\mathcal{D}}_{\mathrm{in}}$ keeps margin $\geq\gamma_{0}$ under the new model $f_{\theta}^{\mathrm{aug}}$.
  2. A2.
    Boundary–sample slack. For every boundary sample $\hat{x}$ we have the (possibly loose) bound $\lvert\gamma(\hat{x})\rvert\leq\delta$ with a fixed constant $\delta>0$.
  3. A3.
    Conservative ordering. We choose $\delta\!\geq\!\gamma_{0}$. This turns A2. into a trivial worst-case cap (it can never be violated) and lets the algebra below produce a bona-fide lower bound—that is, the right-hand side never exceeds the true minimum margin.
Two exhaustive cases.
Let $m_{\mathrm{bd}}:=|\widehat{\mathcal{D}}_{\mathrm{bd}}|$ and $m_{\mathrm{in}}:=|\widehat{\mathcal{D}}_{\mathrm{in}}|$.
Case 1: $\boldsymbol{m_{\mathrm{bd}}=0}$ (BCR $=0$). The augmented set contains only $\mathcal{D}\cup\widehat{\mathcal{D}}_{\mathrm{in}}$. By Assumption A1. every margin is still $\geq\gamma_{0}$, hence $\gamma_{\min}(\tilde{\mathcal{D}})\geq\gamma_{0}.$ The bound in the statement reduces to $\gamma_{0}-\delta(1-0)=\gamma_{0},$ so the inequality is tight in this corner case.
Case 2: $\boldsymbol{m_{\mathrm{bd}}>0}$ (BCR $>0$). Now at least one boundary sample is present. By A2. its margin is $\geq-\delta$ and, because $\delta\geq\gamma_{0}$ (A3.), also $\geq\gamma_{0}-\delta$. Combining with Assumption A1. for all other points gives
$\gamma_{\min}(\tilde{\mathcal{D}})\;\geq\;\min\bigl{\{}\gamma_{0},\;\gamma_{0}-\delta\bigr{\}}\;=\;\gamma_{0}-\delta.$
Because BCR$\in(0,1]$, $\gamma_{0}-\delta(1-\mathrm{BCR})\;\leq\;\gamma_{0}-\delta,$ and the claimed inequality again holds.
Unifying the two cases.
Observe that in Case 1 $(1-\mathrm{BCR})=1$ while in Case 2 $(1-\mathrm{BCR})<1$. Hence the single formula
$\boxed{\;\gamma_{\min}\!\bigl{(}\tilde{\mathcal{D}}\bigr{)}\;\;\geq\;\;\gamma_{0}-\delta\bigl{(}1-\mathrm{BCR}\bigr{)}\;}$
is simultaneously valid for both scenarios, completing the proof.
F.2 Proof of Theorem 4.5 (On-Manifold Vicinal Risk Theorem)
Notation recap.
For each original training pair $(t_{i},c_{i})$:
  • $\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})$ is the vicinal distribution produced by LLM-based interpolation.
  • $\hat{t}\sim\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})$ is one synthetic neighbor.
  • $\phi:\mathcal{T}\!\to\!\mathbb{R}^{d^{\prime}}$ is the frozen encoder; $\mathcal{M}_{c_{i}}=\{\phi(t):y=c_{i}\}$ is the class-$c_{i}$ manifold.
  • $\mathbf{L}(\cdot,\cdot)$ is the loss (cross-entropy, hinge, …).
  • $f_{\theta}$ is the classifier that we are analyzing.
Assumptions for this theorem.
  1. A1.
    Manifold-preservation.  For every $i$ and any neighbor $\hat{t}\sim\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})$,
    $\Pr\!\bigl{[}\phi(\hat{t})\!\notin\!\mathcal{M}_{c_{i}}\bigr{]}\;\leq\;\delta,$
    where $\delta\in(0,1)$ is the constant proved in Theorem 4.2.
  2. A2.
    Uniform bound.  The same $\delta$ works for all training indices $i$ (if they differ, use the worst-case $\delta=\max_{i}\delta_{i}$).
  3. A3.
    Finite loss.$\mathbf{L}(f_{\theta}(\phi(\hat{t})),c_{i})$ is integrable so that all expectations are well defined. [^2] A bounded loss (e.g. cross-entropy with label smoothing) or a sub-Gaussian surrogate satisfies this automatically.
Vicinal risk decomposition.
The VRM objective is
$R_{vrm}(f_{\theta})=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\hat{t}\sim\mathcal{V}_{\mathcal{L}_{\tau}}(t_{i})}\bigl{[}\mathbf{L}(f_{\theta}(\phi(\hat{t})),c_{i})\bigr{]}.$
Fix $i$ temporarily. Apply A1. and the law of total expectation [41]:
$\mathbb{E}_{\hat{t}}\bigl{[}\mathbf{L}(\cdot)\bigr{]}=(1-\delta)\,\mathbb{E}\!\bigl{[}\mathbf{L}(\cdot)\mid\phi(\hat{t})\!\in\!\mathcal{M}_{c_{i}}\bigr{]}+\;\delta\,\mathbb{E}\!\bigl{[}\mathbf{L}(\cdot)\mid\phi(\hat{t})\!\notin\!\mathcal{M}_{c_{i}}\bigr{]}.$
Define
$R_{\text{on}}:=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\!\bigl{[}\mathbf{L}(\cdot)\mid\phi(\hat{t})\in\mathcal{M}_{c_{i}}\bigr{]},\qquad R_{\text{off}}:=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\!\bigl{[}\mathbf{L}(\cdot)\mid\phi(\hat{t})\notin\mathcal{M}_{c_{i}}\bigr{]}.$
Summing the displayed equality over $i=1,\dots,n$ and using A2. yields
$\boxed{\;R_{vrm}(f_{\theta})=(1-\delta)\,R_{\text{on}}+\delta\,R_{\text{off}}\;}$
with $0\leq\delta\ll 1$ coming from Theorem 4.2. Because $(1-\delta)$ is the probability of drawing an on-manifold neighbor, the VRM risk is evaluated on-manifold with probability at least $1-\delta$, completing the proof.
F.3 Proof of Theorem 4.6 (Boundary-Coverage $\Rightarrow$ Vicinal-Risk Reduction)
Notation recap.
  • $\tilde{\mathcal{D}}=\mathcal{D}\cup\widehat{\mathcal{D}}$ is the augmented training set; $\widehat{\mathcal{D}}$ contains $m$ synthetic points.
  • $\widehat{\mathcal{D}}_{\mathrm{bd}}$ / $\widehat{\mathcal{D}}_{\mathrm{in}}$ synthetic boundary / interior subsets; $\mathrm{BCR}:=|\widehat{\mathcal{D}}_{\mathrm{bd}}|/m$.
  • $\gamma(\hat{x})$ logit margin of $\hat{x}$; $\gamma_{0}:=\min_{x\in\mathcal{D}}\gamma(x)$.
  • $\delta$ upper bound on $|\gamma(\hat{x})|$ for $\hat{x}\in\widehat{\mathcal{D}}_{\mathrm{bd}}$ (Theorem 4.4).
  • $\mathbf{L}$ supervised loss and $L$ is the Lipschitz constant of $\mathbf{L}$ in its first (logit) argument.
  • $f_{\theta}^{\mathrm{orig}}$ model trained on $\mathcal{D}$; $f_{\theta}^{\mathrm{aug}}$ model after VRM training on $\tilde{\mathcal{D}}$.
Assumptions used in the proof.
  1. A1.
    Lipschitz loss. $\mathbf{L}(z_{1},y)-\mathbf{L}(z_{2},y)\leq L\|z_{1}-z_{2}\|_{2}$  [42].
  2. A2.
    Non-decreasing margins. VRM retraining never reduces the margin of a synthetic point: $\gamma_{\,\mathrm{aug}}(\hat{x})\geq\gamma_{\,\mathrm{orig}}(\hat{x})$.
  3. A3.
    Boundary slack. $|\gamma(\hat{x})|\leq\delta$ for all $\hat{x}\in\widehat{\mathcal{D}}_{\mathrm{bd}}$.
  4. A4.
    Interior margin. $\gamma(\hat{x})\geq\gamma_{0}$ for all $\hat{x}\in\widehat{\mathcal{D}}_{\mathrm{in}}$.
(Last three assumptions are justified in Appendix F.1.)
Pointwise loss drop.
For any synthetic $(\hat{x},\hat{y})$, combine A1. + A2. to obtain
$\mathbf{L}\bigl{(}f_{\theta}^{\mathrm{aug}}(\hat{x}),\hat{y}\bigr{)}-\mathbf{L}\bigl{(}f_{\theta}^{\mathrm{orig}}(\hat{x}),\hat{y}\bigr{)}\;\leq\;-L\,\gamma(\hat{x}).$
Separate boundary vs. interior sums.
Sum (F.3) over the two synthetic subsets and divide by $m$:
$R_{vrm}\!\bigl{(}f_{\theta}^{\mathrm{aug}}\bigr{)}-R_{vrm}\!\bigl{(}f_{\theta}^{\mathrm{orig}}\bigr{)}$ $\leq-\frac{L}{m}\sum_{\hat{x}\in\widehat{\mathcal{D}}_{\mathrm{bd}}}\!\!\gamma(\hat{x})-\frac{L}{m}\sum_{\hat{x}\in\widehat{\mathcal{D}}_{\mathrm{in}}}\!\!\gamma(\hat{x})$
A3.,A4.${\leq}-L\,\delta\,\frac{|\widehat{\mathcal{D}}_{\mathrm{bd}}|}{m}\;-\;L\,\gamma_{0}\,\frac{|\widehat{\mathcal{D}}_{\mathrm{in}}|}{m}.$
Recognize the fractions as $\mathrm{BCR}$ and $1-\mathrm{BCR}$:
$R_{vrm}\!\bigl{(}f_{\theta}^{\mathrm{aug}}\bigr{)}-R_{vrm}\!\bigl{(}f_{\theta}^{\mathrm{orig}}\bigr{)}\;\leq\;-L\,\gamma_{0}\,\mathrm{BCR}\;+\;L\,\delta\,(1-\mathrm{BCR}).$
Express the bound in compact form.
Choose $\eta=\delta/\gamma_{0}\;(0<\eta<1)$ and rewrite (F.3) as
$R_{vrm}\!\bigl{(}f_{\theta}^{\mathrm{aug}}\bigr{)}-R_{vrm}\!\bigl{(}f_{\theta}^{\mathrm{orig}}\bigr{)}\;\leq\;-L\,\gamma_{0}\,\bigl{[}\mathrm{BCR}-\eta\bigr{]}\;+\;\mathcal{O}(\delta),$
where the $\mathcal{O}(\delta)$ term hides constants independent of $\mathrm{BCR}$ when $\delta$ is small. Hence whenever $\mathrm{BCR}>\eta=\delta/\gamma_{0}$ (the typical case in our experiments), the vicinal risk strictly decreases after VRM training.
F.4 Proof of Theorem 4.8 (Pulling Well-Aligned Nodes)
Notation.
  • $h^{(\ell)}_{v}$ – representation of node $v$ after $\ell$ GNN layers.
  • $\mathcal{M}_{c}$ – representation manifold of class $c$ in $\mathbb{R}^{p}$.
  • $d_{\ell}:=\operatorname{dist}\bigl{(}h^{(\ell)}_{\hat{v}},\mathcal{M}_{c}\bigr{)}$ – Euclidean distance of synthetic node $\hat{v}$ to $\mathcal{M}_{c}$.
  • $\mathbf{W}$ – linear map inside the GNN layer; $\alpha\in(0,1)$, $\beta_{\hat{v}u}\geq 0$ with $\sum_{u\in\mathcal{N}^{\prime}(\hat{v})}\beta_{\hat{v}u}=1$.
Layer-wise update.
For node $\hat{v}$ the generic message-passing rule (Equation 1) reads
$h^{(\ell+1)}_{\hat{v}}=\alpha\,\mathbf{W}h^{(\ell)}_{\hat{v}}+(1-\alpha)\sum_{u\in\mathcal{N}^{\prime}(\hat{v})}\beta_{\hat{v}u}\,\mathbf{W}h^{(\ell)}_{u}.$
Assumptions.
  1. A1.
    Non-expansive layer:  spectral norm $\|\mathbf{W}\|_{2}\leq L<1$.
  2. A2.
    Well-aligned neighbors:$\operatorname{dist}\!\bigl{(}h^{(\ell)}_{u},\mathcal{M}_{c}\bigr{)}\leq\varepsilon$ for every $u\in\mathcal{N}^{\prime}(\hat{v})$.
  3. A3.
    Convexity of weights:$\alpha\in(0,1)$ and $\{\beta_{\hat{v}u}\}$ form a convex combination.
(All three conditions are satisfied by common GNNs and our confidence-based edge assignment.)
Contractive bounds after applying $\mathbf{W}$.
By A1., linear mapping $\mathbf{W}$ shrinks distances by at most factor $L$:
[…]
Distance after aggregation.
Because the right-hand side of (U) is a convex combination (A3.), the triangle inequality and (B) give
$d_{\ell+1}:=\operatorname{dist}\bigl{(}h^{(\ell+1)}_{\hat{v}},\mathcal{M}_{c}\bigr{)}\;\leq\;\alpha Ld_{\ell}\;+\;(1-\alpha)L\varepsilon.$
Geometric contraction (when $d_{\ell}\!\geq\!\varepsilon$).
If the current distance satisfies $d_{\ell}\geq\varepsilon$, then $(1-\alpha)L\varepsilon\leq(1-\alpha)Ld_{\ell}$, so (C) becomes
$d_{\ell+1}\leq\underbrace{\bigl{[}\alpha L+(1-\alpha)L\bigr{]}}_{=\;\alpha L}\,d_{\ell}=:\gamma\,d_{\ell},$
with
$\gamma:=\alpha L\quad\text{and}\quad 0<L<1,\;0<\alpha<1\;\Longrightarrow\;\gamma\in(0,1).$
Exponential convergence.
Iterate the inequality whenever $d_{k}\geq\varepsilon$:
$d_{\ell}\;\leq\;\gamma^{\ell-\ell_{0}}\,d_{\ell_{0}}\quad(\ell\!\geq\!\ell_{0}),$
demonstrating exponential decay toward $\mathcal{M}_{c}$. Once the distance falls below $\varepsilon$, neighbors and self-term are already within the same $\varepsilon$-tube, so all subsequent distances stay bounded by $L\varepsilon$.
Appendix G Notation Table
Table 9: Summary of notation used throughout the paper.
Symbol Meaning
$\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$ Attributed graph: nodes $\mathcal{V}$, edges $\mathcal{E}$, texts $\mathcal{T}$
$\mathcal{G^{\prime}}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$ Augmented graph: augmented nodes $\mathcal{V}^{\prime}$, augmented edges $\mathcal{E}^{\prime}$, augmented texts $\mathcal{T}^{\prime}$
$n=|\mathcal{V}|$ Number of nodes
$\mathcal{X}\in\mathbb{R}^{n\times d}$ Node feature matrix ($d$ features per node)
$\mathcal{Y}=\{y_{v}\}$ Node labels, $y_{v}\in\{1,\dots,C\}$
$\mathcal{P}$ Empirical distribution on $(\mathcal{X},\mathcal{Y})$
$f_{\theta}$ Node-level classifier (e.g. GNN)
$\phi:\mathcal{T}\!\to\!\mathbb{R}^{d^{\prime}}$ Frozen text encoder; $z_{v}=\phi(t_{v})$
$\mathcal{M}_{c}$ Representation manifold of class $c$; $\mathcal{M}=\bigcup_{c}\mathcal{M}_{c}$
$\mathcal{C}_{t}$ Set of tail (minority) classes
$\mathcal{N}_{k}(v)$ $k$-hop / $k$-NN neighborhood of node $v$
$\mathcal{N}^{(\text{tail})}_{k}(v)$ Vicinal twins of $v$ (tail nodes in $\mathcal{N}_{k}(v)$)
$\mathcal{L}_{\tau}$ Temperature-controlled LLM generator (temperature $\tau$)
$\hat{t}_{ij}$ Class-consistent text generated from vicinal twins $(v_{i},v_{j})$
$\bar{\mathcal{V}},\bar{\mathcal{T}}$ Sets of synthetic nodes and their texts
$\kappa_{\omega}$ Confidence function $\kappa:[0,1]$ for edge assignment
$\mathcal{N}_{k}(\hat{v})$ Top-$k$ neighbors chosen for synthetic node $\hat{v}$
$\alpha$, $\beta_{vu}$ Self-weight and neighbor weights in GNN aggregator
$h^{(\ell)}_{v}$ Hidden representation of $v$ at layer $\ell$
$\mathbf{W}$ Linear transformation inside a GNN layer
$\gamma(x)$ Logit margin of sample $x$
$\gamma_{0}$ Minimum margin on original training set
$\mathrm{BCR}$ Boundary-Coverage Rate (Def. 4.3)
$R_{vrm}$ Vicinal risk under VRM
$L$ Lipschitz constant of the loss $\mathbf{L}$
$\delta,\eta$ Constants in margin / risk bounds